8. Paxton Corporation uses the weighted-average method in its process costing system. The Molding Department is the second department in its production process. The data below summarize the department's operations in January.

Beginning work in process inventory ........ Transferred in from the prior department during January ................ Complet

The accounting records indicate that the conversion cost that had been assigned to beginning work in process inventory was $10,973 and a total of $268,107 in conversion costs were incurred in the department during January.
What was the cost per equivalent unit for conversion costs for January in the Molding Department? (Round off to three decimal places.)

_________________________________________________________________

9. Barnett Corporation uses the weighted-average method in its process costing system. The company adds materials at the beginning of the process in Department M. Conversion costs were 75% complete with respect to the 4,000 units in work in process at May 1 and 50% complete with respect to the 6,000 units in work in process at May 31. During May, 12,000 units were completed and transferred to the next department. An analysis of the costs relating to work in process at May 1 and to production activity for May follows:

Work in process May 1......... Costs added during May...... Materials $13,800 $42,000 Conversion $3,740 $26,260

The total cost per equivalent unit for May was:
Answer #1

8)

Equivalent units = Completed units + (Ending units X % of completion)

= 50,300 + (9,400 X 20%)

= 52,180

Total conversion cost = $10,973 + $268,107 = $279,080

Cost per equivalent unit for conversion cost = $279,080 / 52,180 = $5.348

9)

Equivalent units:

Material = 12,000 + 6,000 = 18,000

Conversion cost = 12,000 + (6,000 X 50%) = 15,000

Total cost:

Material = $13,800 + $42,000 = $55,800

Conversion cost = $3,740 + $26,260 = $30,000

Cost per equivalent unit:

Material = $55,800 / 18,000 = $3.10

Conversion cost = $30,000 / 15,000 = $2

Total cost per equivalent unit = $3.10 + $2 = $5.10